OpenStudy (anonymous):

Please help! will fan Verify the basic identity. What is the domain of validity? cot theta= cos theta (csc theta)

OpenStudy (anonymous):

@hero ?

OpenStudy (anonymous):

@jim_thompson5910 ?

OpenStudy (tkhunny):

cotangent = cosine / sine cosecant = 1 / sine Think about everywhere except where the sine function is zero.

OpenStudy (anonymous):

hmm, okay. What should i do next?

OpenStudy (tkhunny):

What did you conclude while you were thinking?

OpenStudy (anonymous):

that if i used reciprocal identities cot theta= cot theta that would prove the identity?

OpenStudy (tkhunny):

Are we proving the identity or discussing its Domain of Validity?

OpenStudy (anonymous):

both

OpenStudy (tkhunny):

I see. A little better punctuation would have cleared that up. 1) Verify - Substitute \(csc(\theta) = 1/sin(\theta)\) and you'll essentlally be done with the verification, since the right-hand side will be the definition of the cotangent. 2) Okay, now is the time to think about validity. Does it work for \(\theta = 0\)?

OpenStudy (anonymous):

no, theta can't be zero, right?

OpenStudy (anonymous):

@tkhunny ?

OpenStudy (tkhunny):

How about \(\theta = \pi\)?

OpenStudy (anonymous):

that can work

OpenStudy (tkhunny):

Really? What is \(\cot(\pi)\)?

OpenStudy (anonymous):

I dont know ! D':

OpenStudy (anonymous):

im so confused

OpenStudy (tkhunny):

You need to know that. How did you determine that \(\theta = 0\) was a problem?

OpenStudy (anonymous):

I was told than the denominator couldn't be zero in class

OpenStudy (tkhunny):

Perfect. Where is \(\sin(\theta) = 0\)? \(\theta \in {0 + k\pi}\), where \(\pi\) is any integer. And we have just concluded the discussion on validity. It's not valid where \(\sin(\theta) = 0\).

OpenStudy (anonymous):

so thats 0 degress and 180 degrees?

OpenStudy (tkhunny):

And any integer multiple of \(\pi\).